### v2

# RLAIF-V: Open-Source AI Feedback Leads to Super GPT-4V Trustworthiness

Abstract
Traditional feedback learning for hallucination reduction relies on labor-intensive manual labeling or expensive proprietary models. This leaves the community without foundational knowledge about how to build high-quality feedback with open-source MLLMs. In this work, we introduce RLAIF-V, a novel framework that aligns MLLMs in a fully open-source paradigm. RLAIF-V maximally explores open-source MLLMs from two perspectives, including high-quality feedback data generation for preference learning and self-feedback guidance for inference-time scaling. Extensive experiments on six benchmarks in both automatic and human evaluation show that RLAIF-V substantially enhances the trustworthiness of models at both preference learning and inference time. RLAIF-V 7B reduces object hallucination by 80.7% and overall hallucination by 33.7%. Remarkably, RLAIF-V 12B further reveals the self-alignment potential of open-source MLLMs, where the model can learn from feedback of itself to achieve super GPT-4V trustworthiness.
Figure 1: (a) This work aims to shift the current paradigm of aligning MLLMs with feedback from superior teachers, to align with feedback from peers exhibiting comparable or equal capabilities. (b) Trustworthiness of RLAIF-V compared to other methods. We assess the generative trustworthiness with human evaluation benchmark MHumanEval [66], and evaluate the discriminative trustworthiness with automatic evaluation benchmark AMBER [58].
1 Introduction
Recent advances in multimodal large language models (MLLMs) mark a significant milestone in AI research [4, 10, 33, 35, 34, 65]. These models are trained on large-scale multimodal corpora and possess profound world knowledge, showing remarkable capabilities in tackling diverse multimodal tasks [27, 39, 43]. However, it has been commonly noticed that MLLMs are prone to confidently generating incorrect content that deviates from human preferences [56, 79, 66, 21]. In order to align MLLMs with human preferences, reinforcement learning from human feedback (RLHF) has been widely used and demonstrates substantial results [56, 66]. However, RLHF depends heavily on labor-intensive human annotations and is consequently hard to cover the widespread misalignment between model and human preferences. Recently, reinforcement learning from AI feedback (RLAIF), which uses the preference collected from labeler models as a proxy of human preference, has shown promising potential as an alternative to RLHF [24].
However, current approaches face two challenges: (1) Infeasible labeler requirement. Existing RLAIF methods, demonstrated at the left top of Figure 1, rely on costly proprietary models to distill feedback from [28, 76, 80]. More critically, this paradigm essentially distills the capability of proprietary models to provide a temporary solution for bridging the performance gap. The community consequently lacks knowledge about how to build high-quality feedback using open-source MLLM labelers of comparable capability, as demonstrated at the left bottom of Figure 1. Simply changing the labeler model from a proprietary model to a weaker open-source model leads to unsatisfactory feedback quality due to their limited capacity [6]. (2) Limited inference-time scaling. Inference-time scaling has drawn great attention from the LLM community and shows promising results [54, 45]. Nevertheless, recent works in MLLMs mainly focus on the preference learning stage to utilize high-quality feedback [66, 78, 74] while omitting the importance of feedback in the inference stage. Besides, aimlessly scaling inference computation budget can hardly contribute to the performance, since accurate feedback guidance plays an important role for effective inference-time scaling.
RLAIF-V addresses these challenges through two key innovations: (1) For high-quality feedback generation, we propose a novel deconfounded candidate response generation strategy for better data efficiency and a divide-and-conquer approach for higher pairwise preference accuracy. The deconfounded strategy accurately exposes the genuine trustworthiness difference within response pairs by generating candidate responses from multiple sampling decoding trials under the same condition. Consequently, confounding factors such as the text style are eliminated, and the feedback focuses on the substantial content of responses. The divide-and-conquer approach decomposes the difficult response-evaluation task into simpler claim-evaluation, which substantially simplifies the task and thus reduces the capacity requirement of labeler models. (2) For inference-time scaling guidance, we propose a novel self-feedback approach based on models aligned with direct preference optimization [48] (DPO). Specifically, we leverage the reward score generated by aligned models as feedback for itself. However, previous works have shown that direct feedback from DPO-aligned models can be biased towards shorter responses due to its objective formulation [49]. We devise a length-normalization strategy to aggregate the token-level scores of each response for bias suppression. Moreover, we also extensively explore the inference-time scaling [54, 45] potential of our RLAIF-V reward on other open-source models, and demonstrate that a single reward model can well generalize to improve the trustworthiness of multiple MLLMs.
Comprehensive experiments on six benchmarks show that RLAIF-V can substantially enhance model trustworthiness without any human or proprietary model intervention. Using feedback from LLaVA-NeXT 34B [35], RLAIF-V 7B significantly reduces the object hallucination on Object HalBench [66] by 80.7%, even surpassing the labeler model. Pushing the limit to an extreme scenario where no stronger models are available, we align OmniLMM 12B [46] with itself as the labeler. Experimental results show that RLAIF-V 12B reduces object hallucination by 76.8% in Object HalBench and overall hallucination by 32.4% in MHumanEval, surpassing GPT-4V by a large margin and revealing the self-alignment potential of open-source MLLMs.
Figure 2: Overview of the RLAIF-V framework. (1) Given the input image and prompt, multiple candidate responses are generated with the deconfounded strategy. (2) Each response is split into atomic claims and assigned with trustworthiness scores separately by an open-source MLLM. (3) During preference learning, the model is aligned under an iterative feedback learning approach which periodically updates the feedback. (4) During inference, performance of the aligned model is further improved by self-feedback guidance.
The contribution of this work can be summarized as fourfold: (1) We present RLAIF-V, a novel framework that aligns MLLMs with open-source feedback. (2) We propose a novel deconfounded and divide-and-conquer approach to generate human-level quality feedback with open-source models. (3) We propose a novel self-feedback guidance for inference-time scaling using aligned MLLMs and devise a simple length-normalization strategy tackling the bias towards shorter responses. (4) We conduct comprehensive experiments to demonstrate the effectiveness of the proposed framework, achieving state-of-the-art performance in trustworthiness among both open-source and proprietary MLLMs. All codes, data, and model weights will be released to facilitate future research.
2 RLAIF-V
In this section, we first elaborate on how to collect high-quality AI feedback from open-source MLLMs by introducing the response generation and feedback annotation process. Then we introduce the iterative feedback learning stage and the self-feedback guidance for inference-time scaling. An overview of the RLAIF-V framework is demonstrated in Figure 2.
2.1 Response Generation
The feedback collected for preference learning is in the form of comparison pairs, where each pair includes a preferred response $y_{w}$ and an inferior response $y_{l}$ to the same input $x$ (including the image and prompt). During training, the model learns preferences by distinguishing the differences between $y_{w}$ and $y_{l}$. However, these differences can be complex and consist of many factors including not only the meaning of content but also textual styles such as the use of specific words or structure of the text, making the learning more difficult.
To expose the genuine differences in trustworthiness between responses, we propose a novel deconfounded strategy to generate candidate responses. Specifically, we ask the model to generate $n$ candidate responses $\{y_{1},y_{2},\cdots,y_{n}\}$ through sampling decoding with different random seeds, where input $x$ and decoding parameters are invariant. In this way, $y_{w}$ and $y_{l}$ are sampled from the same distribution and consequently share similar textual styles and linguistic patterns. During training, the model can effectively concentrate on the differences in trustworthiness. In our experiments, we find the deconfounded strategy can significantly improve the learning efficiency (see Section 3.4).
2.2 Feedback Annotation
Evaluating the quality of model responses is a challenging task even for human annotators due to the complexity of full responses. Existing methods using models as labelers rely on costly API of proprietary models with extraordinary instruction-following and task-solving capabilities [69], resulting in scalability issues. In contrast, we employ a divide-and-conquer approach to simplify the task to achieve more reliable results from open-source MLLMs. The detail of collecting high-quality feedback with this approach is described as follows:
Divide. The complexity of full responses makes the holistic assessment of response quality hard to acquire based on existing open-source MLLMs [6]. One of the important complexity causes is that a full response might contain multiple statements and specific textual structure which interferes with the recognition of incorrect spans. To make such a complicated task solvable, we decompose the response evaluation into atomic claim evaluation, as shown in Figure 2. Specifically, we prompt a large language model to split a response $y$ into atomic claims $\{c_{1},c_{2},\cdots,c_{m}\}$, which can be evaluated separately, by extracting facts excluding opinions and subjective statements.
Conquer. To access the trustworthiness quality of a claim $c$ (e.g., “The clock reads around 11:20.”), we first convert it into a polar question like “Does the clock read around 11:20?”, which can be answered with simply yes or no, without introducing any extra content. For each atomic polar question, we ask an open-source MLLM to generate the confidence of agreement and disagreement as the claim score $s_{c}=(p_{yes},p_{no})$, where $p_{yes}$ is the probability of answering with “Yes” or “yes” and $p_{no}$ is the probability of answering with “No” or “no”. A higher $p_{yes}$ score suggests the corresponding claim is considered more trustworthy by the labeler model. The scores collected in this way are generally more accurate compared with directly querying the evaluation result of the full response since the claims are simpler in both structure and content.
Combine. After obtaining the quality assessment of each claim, we finally combine them into the score of the whole response. For each response, we denote the number of claims having $p_{no}>p_{yes}$ as $n_{rej}$, measuring how many incorrect claims are recognized by the labeler model. We use $-n_{rej}$ as the final score $S$ of the response, where a higher score indicates less incorrectness of the content. Given the score of each response, we can now construct a preference dataset for training. For each instruction $x$, we keep all response pairs $(y,y^{\prime})$ such that $S>S^{\prime}$ and choose the higher score response $y$ as the preferred response. To save the training cost, we randomly sample at most 2 pairs for each instruction and we find such a filtering process only causes minor performance drops. To prevent the potential length bias, we drop pairs in which $y_{w}$ is too short before training to ensure the average length difference of $y_{w}$ and $y_{l}$ is less than one word.
2.3 Iterative Feedback Learning
DPO is widely used to align MLLMs with human preferences. However, naive DPO faces the distribution shift problem, i.e., the preference data is static during the training process while model output distribution is constantly shifting [14]. As a result, the data distribution might deviate from the expected feedback distribution and cause suboptimal alignment results.
We follow [17] to train the model iteratively. Specifically, we select $N$ multimodal instructions at the beginning of each iteration and leverage the deconfounded strategy to generate $n$ candidate responses for each instruction with the latest instruction model $M_{i}$. We assign each response with a trustworthiness score through the divide-and-conquer approach using the labeler model $L$ and construct comparison pairs $D_{i}$ for training. Then we train the $M_{i}$ with direct preference optimization on $D_{i}$ to get $M_{i+1}$, which is used as the instruction model of the next iteration. In this way, the feedback distribution can be updated in an iterative manner, resulting in better learning efficiency.
2.4 Self-Feedback for Inference-time Scaling
After iterative learning on diverse high-quality feedback, the MLLM itself is not only a trustworthy policy model but also a reward function by the optimization objective of DPO [48] and the reward is formulated as:
$\begin{split}r(y)&=\beta\log{\frac{\pi_{\theta}(y)}{\pi_{\text{ref}}(y)}}=\beta\sum_{t}^{T}{\log{\frac{\pi_{\theta}(y_{t}|y_{<t})}{\pi_{\text{ref}}(y_{t}|y_{<t})}}},\end{split}$ (1)
where $\beta$ is a parameter controlling the deviation from the base reference policy $\pi_{ref}$, $y$ is the response token sequence, $T$ is the length or response and $\pi_{\theta}$ is the model after DPO training. We hide the prompt condition $x$ in equations for simplicity. Previous works have shown that DPO-aligned reward $r(y)$ can be biased towards shorter responses due to its objective formulation [49]. We tackle this bias by averaging all token-level scores to get the final response score $r(y)=\frac{\beta}{T}\log{\frac{\pi_{\theta}(y)}{\pi_{\text{ref}}(y)}}$.
We then use the normalized reward as self-feedback guidance for inference-time scaling. Specifically, we follow [54] to perform best-of-N (BoN) selection based on multiple sampled responses of the same prompt. Specifically, we choose the response with the highest score among the N candidate responses as the model prediction. To amplify the candidate response diversity, we follow existing works [54] to apply commonly used nucleus sampling [19] for decoding.
3 Experiments
In this section, we empirically investigate the effectiveness of RLAIF-V in aligning MLLMs through open-source feedback. In addition to evaluating model performance regarding trustworthiness and helpfulness, we also analyze the efficacy of different components, the compatibility with other methods, and the generalizability of feedback data collected with RLAIF-V.
Model Size Feedback Object MHum. MMHal- AMBER MM- RefoMB
HalBench Bench Star
Rsp. $\downarrow$ Men. $\downarrow$ Rsp. $\downarrow$ Score Hall. $\downarrow$ Acc. F1 Avg. Trust. Win.
VCD [25]
(CVPR’24)
7B 48.8 24.3 67.1 2.12 54.2 71.8 74.9 33.8 39.9 16.7
Less-is-more [71]
(ACL’24)
7B 40.3 17.8 63.7 2.33 50.0 72.4 75.8 32.9 51.1 16.2
OPERA [21]
(CVPR’24)
7B 45.1 22.3 63.0 2.15 54.2 75.2 78.3 32.9 33.8 13.1
CCA-LLaVA [64]
(NeurIPS’24)
7B 46.7 23.8 68.5 1.92 61.5 77.7 81.9 32.1 41.9 21.7
Qwen-VL-Chat [4]
(arXiv’23)
10B 40.4 20.7 61.0 2.76 38.5 81.9 86.4 34.5 40.9 17.7
LLaVA-NeXT [35]
(arXiv’24)
34B 12.6 6.4 53.4 3.31 34.4 81.4 85.4 51.6 44.4 35.4
MiniGemini [31]
(arXiv’24)
34B 14.5 8.0 59.6 3.08 38.5 82.6 87.6 45.5 50.0 36.9
HA-DPO [76]
(arXiv’23)
7B Rule 39.9 19.9 53.4 1.98 60.4 75.2 79.9 32.9 39.9 17.2
POVID [78]
(arXiv’24)
7B Rule 48.1 24.4 67.8 2.08 56.2 82.9 87.4 34.3 44.4 13.6
LLaVA-RLHF [56]
(arXiv’23)
13B Human 38.1 18.9 72.6 2.02 62.5 79.7 83.9 34.2 26.3 17.2
Silkie [28]
(EMNLP’24)
10B GPT-4V 27.1 13.4 54.1 3.19 32.3 82.2 87.6 33.6 38.9 21.2
RLHF-V [66]
(CVPR’24)
13B Human 12.2 7.5 55.5 2.45 51.0 72.6 75.0 33.2 41.4 17.7
AMP-MEG [74]
(NeurIPS’24)
13B Rule 31.7 20.6 54.8 3.08 36.5 79.5 84.6 34.8 30.3 14.6
LLaVA 1.5 7B 54.5 27.8 67.1 1.86 63.5 73.5 77.7 33.3 36.9 16.2
+ RLAIF-V 7B LLaVA-NeXT 10.5
+44.0
5.2
+22.6
44.5
+20.6
2.95
+1.1
32.3
+31.2
76.8
+3.3
84.5
+6.8
35.4
+2.1
47.2
+10.3
22.5
+6.3
+ RLAIF-V BoN 7B  LLaVA-NeXT 6.8
+3.7
3.8
+1.4
39.7
+4.8
3.07
+0.1
28.1
+4.2
N/A N/A N/A 55.7
+8.5
24.4
+1.9
OmniLMM 12B 19.4 10.9 52.7 3.06 36.5 86.5 89.5 39.7 44.7 18.5
+ RLAIF-V 12B self 4.5
+14.9
2.9
+8.0
35.6
+17.1
3.15
+0.1
32.3
+4.2
88.0 +1.5 90.9 +1.4 40.9
+1.2
58.1
+13.4
28.3
+9.8
+ RLAIF-V BoN 12B self 4.5 +0.0 2.6 +0.3 29.5
+6.1
3.44
+0.3
26.0 +6.3 N/A N/A N/A 62.9 +4.8 30.3
+2.0
GPT-4V [43] - Unknown 13.6 7.3 45.9 3.49 28.1 83.4 87.4 50.4 50.0 50.0
Table 1: Main experimental results. We report hallucination rates in different granularities including response-level (Rsp.) and mention-level (Men.). MHum.: MHumanEval, Hall.: Hallucination Rate, Trust.: trustworthiness win rate, Win.: overall win-rate. The best results are shown in bold. BoN: apply RLAIF-V 7B and RLAIF-V 12B self-feedback for best-of-N, we sample 32 and 16 samples respectively to control evaluation cost. N/A: multi-choice and yes-no question do not have BoN results since these questions only requires single token.
3.1 Experimental Setup
We introduce models, training data, evaluation benchmarks, baselines, and other implementation details. All experiments are conducted based on LLaVA 1.5 7B [33] unless otherwise specified.
Models. We present two settings to align MLLMs with the RLAIF-V framework. First, we use LLaVA 1.5 [33] as the instruction model and LLaVA-NeXT [35] as the labeler model, demonstrating the effectiveness of open-source feedback. Second, we use OmniLMM [46] as both the instruction model and labeler model, representing the extreme scenario where no stronger models are available.
Training Data. The diversity of instructions can be critical for models to learn generalizable preferences. In practice, we use instructions collected from a diverse range of datasets, including MSCOCO [32], ShareGPT-4V [7], MovieNet [20], Google Landmark v2 [61], VQA v2 [15], OKVQA [40], and TextVQA [53]. In addition, we adopt image description prompts introduced in [66] to construct long-form image describing instructions.
Data ObjHal. AMBER
Rsp. $\downarrow$ Men. $\downarrow$ Acc. F1
RLHF-V [66] 28.5 12.3 76.4 84.6
RLAIF-V 10.1 4.7 80.1 86.1
w/o deconfounding 25.7 11.8 73.3 83.0
Table 2: Experimental results of different response generation methods. ObjHal.: Object HalBench.
Evaluation. We evaluate models from two perspectives, including trustworthiness reflecting the hallucination degree, and helpfulness reflecting the general capability. For trustworthiness, we perform evaluation on five benchmarks:
(1) Object HalBench [51] is a widely adopted benchmark for assessing common object hallucination in detailed image descriptions. We follow [66] to use 8 diverse prompts to improve the evaluation stability. We report the response-level hallucination rate (i.e., the percentage of hallucinated responses) and the mention-level hallucination rate (i.e., the percentage of hallucinated objects).
(2) MMHal-Bench [56] evaluates response-level hallucination rate and informativeness. It asks GPT-4 [44] to compare model outputs with human responses and object labels for evaluation.
(3) MHumanEval [66] comprises 146 samples collected from both Object HalBench (50) and MMHal-Bench (96) to provide a more comprehensive evaluation over both long-form description and short-form questions. We only label the response-level hallucination rate to control the cost.
(4) AMBER [58] is a multi-dimensional hallucination benchmark comprising more than 15k samples. We use the discriminative part and report the accuracy and F1 metric.
The above trustworthiness evaluations are either limited to common object hallucination, which is mostly eliminated, constrained format (e.g., yes-no choices) or manual labeling. To reliably and automatically assess the trustworthiness of MLLMs under any format, we construct a novel Reliable Free-format Multimodal Benchmark (RefoMB) containing 120 images and 360 instructions covering 8 critical tasks such as mechanical reasoning [38] and image perception [2]. Following [34], we assess the performance of MLLMs by comparing the model response with GPT-4V response regarding both trustworthiness and helpfulness. We calculate the trustworthiness win rate and overall win rate based on the evaluation review. Each instruction is paired with a thoroughly written image description as the reference, achieving a notable 96% human agreement. Results on the dev split (99 instructions) are reported in this section to save evaluation costs, we present more details and the test split (261 instructions) results in the Appendix.
Data Agree. ObjHal. AMBER
Rsp. $\downarrow$ Men. $\downarrow$ Acc. F1
VL-Feedback [28] 92.3% 37.9 21.0 72.8 82.6
RLAIF-V 96.7% 20.6 10.4 80.5 86.0
     w/o d&c 66.7% 53.1 26.2 73.5 77.7
     w/ smaller labeler 90.0% 33.0 17.5 75.2 78.2
Table 3: Performance comparison of different feedback collection methods. We conduct the experiment on various labeler models. ObjHal.: Object HalBench. smaller labeler: OmniLMM 12B, Agree.: Human agreement of the constructed pairs, d&c: divide-and-conquer strategy. VL-Feedback collects high-quality feedback from GPT-4V.
For helpfulness, we adopt MMStar [8], which is a comprehensive benchmark containing 1500 challenge samples collected from 6 popular multimodal benchmarks [70, 36, 37, 26, 38, 23], covering 6 core capabilities and 18 detailed axes. We report the overall score on this benchmark.
Baselines. We compare our model with state-of-the-art baselines of different types, including general baselines with strong performance, baselines trained with feedback data, baselines reduce hallucination without feedback data and proprietary baselines.
(1) General baselines. We adopt LLaVA 1.5 [33], Qwen-VL-Chat [4], OmniLMM [46], LLaVA-NeXT [35], MiniGemini [31] as representative general baselines.
(2) Baselines tailored for feedback learning. RLHF-V [66] collects fine-grained correctional human feedback and trains the model with dense direction preference optimization. Silkie [28] utilizes GPT-4V to collect feedback. POVID [78] and AMP-MEG [74] apply heuristic rules to pair responses generated under difference condition.
(3) Baselines tailored for hallucination reduction without feedback. VCD [25] contrasts model logits derived from original and distorted visual input to reduce the over-reliance on statistical bias and unimodal priors. OPERA [21] introduces a penalty term on the model logits. Less-is-more [71] proposes a selective end-of-sentence (EOS) special token supervision loss and data filtering strategy. CCA-LLaVA [64] mitigates hallucination by applying a novel concentric causal attention.
(4) Proprietary baseline. We also include GPT-4V [43] as strong reference to evaluate the gap between the open-source models and proprietary models.
Implementation Details. We use the Nous-Hermes-2-Yi-34B [3] version of LLaVA-NeXT and the no-RLHF version of OmniLMM [46] as labeler models. For each iteration, we train the model with DPO for 4 epochs, with a learning rate 5e-7, beta 0.1, and batch size of 8. We train both RLAIF-V 7B and RLAIF-V 12B for 4 iterations, where we use 4k instructions to collect feedback at each iteration. In summary, it costs 48h and 50h for data collection of 7B and 12B models, and costs 6h and 8h for training separately, using an 8xA100 80G machine. For the best-of-N setting, we sample 32 and 16 candidate responses for RLAIF-V 7B and RLAIF-V 12B respectively to control the evaluation cost.
3.2 Main Results
The main experimental results are reported in Table 1, from which we observe that: (1) RLAIF-V achieves state-of-the-art performance in trustworthiness among open-source models and even surpasses proprietary models such as GPT-4V. The framework significantly reduces the object hallucination rate of LLaVA 1.5 and OmniLMM by 80.7% and 76.8% relative points on Object HalBench. For the overall hallucination rate, RLAIF-V 12B achieves 35.6% on MHumanEval, surpassing GPT-4V by a large margin. The reduction of hallucination is consistent among multiple benchmarks including MMHal-Bench, AMBER, and RefoMB. (2) RLAIF-V achieves promising performance in response helpfulness, where the results on MMStar are improved compared to the base model. This shows that RLAIF-V can enhance the trustworthiness of MLLMs without sacrificing the performance of other tasks. (3) Using OmniLMM as both the instruction model and the labeler model, RLAIF-V 12B achieves significant hallucination reduction on multiple benchmarks and comparable helpfulness. Remarkably, RLAIF-V 12B outperforms GPT-4V in trustworthiness on Object HalBench, MHumanEval, AMBER, and RefoMB, by substantial margins. The results demonstrate a promising path to achieve self-alignment of leading-edge MLLMs. (4) Self-feedback guidance improves the trustworthiness of both RLAIF-V 7B and RLAIF-V 12B on multiple benchmarks with best-of-N selection, demonstrating the effectiveness of RLAIF-V reward at inference-time.
3.3 Ablation Study
To investigate the contribution of different components in RLAIF-V, we perform an ablation study.
Ablation of response generation approach. To quantify the advantage of the deconfounded candidate response generation strategy, we conduct an experiment based on the RLHF-V dataset [66]. We compare the performance of model trained under three settings: (1) RLHF-V, the model is directly aligned with human feedback data; (2) RLAIF-V, we collect high-quality feedback from LLaVA-NeXT based on original multimodal instructions in RLHF-V dataset using the RLAIF-V framework; (3) RLAIF-V w/o deconfounding, we replace the preferred responses generated under the deconfounded strategy with original human annotations.
From experimental results in Table 2, we observe that the model trained with our deconfounded responses achieves the best performance on both tasks. Changing preferred responses with high-quality human annotated responses, though improving the feedback precision and response quality, exhibits significant performance loss. We hypothesize this action introduces more non-robust shallow patterns into the training data and thus harms the learning efficiency. Moreover, the performance of our method even surpasses training on human-annotated correctional feedback by a large margin. After analyzing the composition of the RLHF-V dataset, we find it only includes a limited selection of models [66] which share limited hallucination distribution similarity with LLaVA 1.5 7B. As a result, the effectiveness of the dataset is significantly diminished. We argue this phenomenon further enhances the importance of the RLAIF-V framework which can efficiently generate high-quality feedback data for any MLLM. We list more details about the hallucination distribution and RLHF-V dataset composition in the Appendix D.
Effect of divide-and-conquer strategy. We compare our divide-and-conquer strategy with direct self-rewarding [69] by replacing only the implementation of response evaluation process. Specifically, self-rewarding asks the labeler model to generate an overall quality score of each candidate response with a long prompt introducing multiple criteria. We assess the human agreement of generated response pairs by the ratio of $(y_{w},y_{l})$ which evaluators agree that $y_{w}\geqslant y_{l}$. Based on results in Table 3, we observe that simply asking open-source models to generate an overall assessment of responses yields unsatisfactory results due to poor feedback quality. In contrast, our method with the divide-and-conquer strategy significantly improves the feedback quality and overall performance on both discriminative and generative tasks. Moreover, we also compare RLAIF-V feedback data with VL-Feedback [28] which collects high-quality feedback from GPT-4V. Results show RLAIF-V achieves higher data quality with a novel divide-and-conquer strategy and better performance by training on the same amount of data.
Figure 3: Results of combining RLAIF-V with other feedback. We report the response-level no-hallucination rate on Object HalBench for generative trustworthiness.
Figure 4: Hallucination reduction of other MLLMs with data from the first training iteration of RLAIF-V 12B. We report the response-level hallucination rate reduction on Object HalBench for generative hallucination and AMBER error rate reduction for discriminative hallucination.
Figure 5: Inference-time scaling across different models. We report the response level no-hallucination rate on Object HalBench for generative trustworthiness. Reference PPL: using perplexity from OmniLMM.
3.4 Analysis
We conduct analysis on the framework considering the following research questions: (1) Is RLAIF-V compatible with other sources of feedback together? (2) Can feedback data collected for one model with RLAIF-V be adopted to enhance the trustworthiness of other MLLMs? (3) How RLAIF-V reward work for inference-time scaling?
RLAIF-V is complementary with existing feedback collection methods. Besides collecting feedback using models as labelers, multiple existing works generate feedback based on heuristic rules or human annotation. We explore the possibility of combining RLAIF-V with other sources of feedback. Results in Figure 3 show that heuristically collected feedback from HA-DPO [76] and human annotated feedback from RLHF-V can further improve the trustworthiness, indicating RLAIF-V is complementary with other types of feedback.
RLAIF-V produces generalizable high-quality feedback. We train different models with the feedback collected during the first iteration of training RLAIF-V 12B. Specifically, we train LLaVA 1.5 7B [33], LLaVA 1.5 13B [33], MiniCPM-V [46], MiniCPM-V 2 [46]. with direct preference optimization and report the trustworthiness improvement in Figure 4. We observe that data collected from OmniLMM (as both the instruction model and labeler model) with RLAIF-V framework can effectively reduce the hallucination of other MLLMs on different benchmarks. Notably, the improvement can be even more significant compared with the OmniLMM which generates the candidate responses. The results demonstrate that feedback from RLAIF-V is generalizable to improve the trustworthiness of different MLLMs.
RLAIF-V reward continuously improves MLLM trustworthiness at inference-time. We explore the effectiveness of RLAIF-V 12B reward on different open-source MLLMs. As results shown in Figure 5, RLAIF-V reward consistently improves the generative trustworthiness of both LLaVA 1.5 7B and Qwen-VL-Chat. We compare the improvement with two baselines which use the perplexity (PPL) of RLAIF-V 12B model or OmniLMM and observe that RLAIF-V reward achieves significantly better results. We also analyze the average length of best-of-N selected responses compared with naive sampled responses and demonstrate that the simple length-normalization method effectively tackles the bias of preferring shorter responses which might cause significant information loss. Specifically, the average length difference count by words is increased from -7.7 (w/o length-normalization) to +3.9 when using RLAIF-V 12B reward for best-of-64 setting of LLaVA 1.5 7B.
4 Related Works
We introduce the most related background works in this section and refer readers to the Appendix for a more detail review of related works.
Learning from Feedback.
Learning from feedback is one of the core techniques in developing advanced LLMs [9, 68, 57] and MLLMs [66, 56, 28, 78], which aligns the model with human preference. Proximal policy optimization (PPO) [52] is recognized as the major technique to directly align models with human preferences through training a reward model on pairwise comparisons of model responses. Rafael et al. [48] propose direct preference optimization to stabilize the training of PPO and is widely adopted by the community recently. However, most multimodal feedback learning methods only utilize the simplicity and training stability of DPO while omits the important fact that DPO actually trains an optimal reward model. As a result, without explore the effectiveness of the continuous rewards, these methods gets suboptimal outcomes.
Feedback Collection for MLLMs. Feedback quality is one of the most important factors for models to align with human preferences. Early works mainly collect high-quality feedback through human labelers which is costly and limited compared with the widespread misalignment problem [66, 56, 16]. To this end, collecting feedback from AI serves an alternative to get rid of human intervention and provides a promising way to guide super-intelligent models [9]. However, existing methods simply distill feedback for MLLMs from proprietary models like GPT-4V, which rely on the superiority of proprietary model over the student model which uses the feedback to improve itself [28]. The concurrent HSA-DPO [63] asks GPT-4 [44] and GPT-4V [43] to detect hallucination from 6k image descriptions. FGAIF [22] asks ChatGPT to split the response into sub-sentences and classify them into either object-existence or attribute or relation relevant. These approaches still depend on strong proprietary models and only tackle MLLM hallucination on the image captioning task regarding three kinds of object-related hallucination. RLAIF-V, on the other hand, strengthens MLLMs with feedback on a diverse range of tasks (e.g., visual question answering [40], scene text understanding [53] and image captioning [32]) under a fully open-source setting. HA-DPO [76], POVID [78], AMP [74] and BPO [47] heuristically construct comparison pairs by either distorting the image, editing the model response or pairing models with different performance.
[…]
5 Conclusion
Aligning models with human preference to reduce MLLM hallucination is a critical target. In this work, we present RLAIF-V, a novel framework that enhances the trustworthiness of MLLMs through open-source AI feedback. Comprehensive experimental results show that our models achieve state-of-the-art performance in both generative and discriminative trustworthiness. We propose a deconfounded sampling and divide-and-conquer strategy to improve the efficiency and quality of feedback. By aligning the model with such high-quality feedback, the trustworthiness can be substantially improved without sacrificing performance on other tasks. Moreover, we propose novel self-feedback guidance for inference-time scaling using the aligned model itself and a simple length-normalization strategy to tackle the bias towards shorter responses. We also demonstrate that feedback generated via the RLAIF-V framework is generalizable to different MLLMs. In the future, we will explore collecting more complex feedback from models to improve logical reasoning and complex task-solving capabilities.
Supplementary Material
Appendix A Extended Related Work
Learning from Feedback.
Learning from feedback is one of the core techniques in developing advanced LLMs [9, 68, 57] and MLLMs [66, 56, 28, 78], which aligns the model with human preference. Proximal policy optimization (PPO) [52] is recognized as the major technique to directly align models with human preferences through training a reward model on pairwise comparisons of model responses. Rafael et al. [48] propose direct preference optimization to stabilize the training of PPO and is widely adopted by the community recently. However, DPO relies on a prepared collection of pairwise data, which remains static during training and consequently causes the distribution shift problem. To mitigate such problem, RLAIF-V adopt an iterative training framework to acquire fresh feedback based on output distribution of current model and use the feedback to update the model.
Feedback Collection for MLLMs. Feedback quality is one of the most important factors for models to align with human preferences. Early works mainly collect high-quality feedback through human labelers which is costly and limited compared with the widespread misalignment problem [66, 56]. To this end, collecting feedback from AI serves an alternative to get rid of human intervention and provides a promising way to guide super-intelligent models beyond human performance [9]. However, existing methods simply distill feedback for MLLMs from proprietary models like GPT-4V, which rely on the superiority of proprietary model over the student model which uses the feedback to improve itself [28]. The concurrent HSA-DPO [63] asks GPT-4 [44] and GPT-4V [43] to detect hallucination from 6k image descriptions and use the output to train a 40B detector model for hallucination detection. It then applies a 34B re-writer model to re-write hallucinated sentences to form preference pairs. FGAIF [22] asks ChatGPT to split the response into sub-sentences and classify them into either object-existence or attribute or relation relevant which are further used to collect feedback from the LLaVA 1.5 13B to get a score of each response. These approaches still depend on strong proprietary models and only tackle MLLM hallucination on the image captioning task regarding three kinds of object-related hallucination. RLAIF-V, on the other hand, strengthens MLLMs with feedback on a diverse range of tasks (e.g., visual question answering [40], scene text understanding [53] and image captioning [32]) under a fully open-source setting. HA-DPO [76], POVID [78] and BPO [47] heuristically construct comparison pairs by either distorting the image or editing the model response. FDPO [16] employs human annotators to collect span-level fine-grained feedback to reduce the hallucination of MLLMs.
[…]
Appendix B RefoMB
In this section, we introduce details about RefoMB and conduct more analyses on it. The benchmark contains 120 images, each annotated with 3 instructions, and assesses 8 core capabilities covering both perception and reasoning.
B.1 GPT-4 as Evaluator
Evaluating the quality of open-ended responses in terms of trustworthiness and helpfulness presents significant challenges. Inspired by the progress of utilizing LLMs to evaluate language models, recent MLLM benchmarks including LLaVA Bench [34] and MMHal-Bench [56] adopt GPT-4 as evaluator to handle the complexity of open-ended responses. However, these benchmarks exhibit divergence from human judgment due to the incompleteness of image information provided to the GPT-4 evaluator, which hinders the reliability of their evaluation results. To address this problem, we propose to annotate each image with a comprehensive description, conveying most of the content in the image. The annotation process of these descriptions is elaborated in the next section. Specifically, the thorough image description each contains 706 words on average. In line with the widely used LLM evaluation benchmark AlpacaEval [29], we utilize GPT-4 to assess response quality by comparing it to the response from a competitor model.
During the evaluation, we pass the comprehensive image description, instruction, and two responses (i.e., from both the model being evaluated and the competitor model) to GPT-4 with the prompt shown in Figure 7. The evaluation focuses on the trustworthiness and overall helpfulness of the responses, where trustworthiness is gauged by the number of hallucinations in the response, and helpfulness is measured by the effectiveness in assisting the user in achieving their goals (i.e., the instruction). With the comprehensive description that encapsulates most content of the image, GPT-4 can follow the aforementioned evaluation criteria more reliably. We select GPT-4V [43] as the competitor model since it is one of the most powerful MLLMs.
Categories Fine-grained Coarse Creative OCR Relation Attribute Logical Time series All
Perception Perception Generation Reasoning Reasoning Reasoning Reasoning
Dev 24 19 12 11 14 12 3 4 99
Test 61 50 33 28 38 32 8 11 261
Total 85 69 45 39 52 44 11 15 360
Table 4: The number of instructions in each category of RefoMB.
Figure 6: RefoMB instructions distribution.
B.2 Benchmark Construction
The construction process of RefoMB involves the image collection and comprehensive description annotation described in B.2.1, as well as the instruction design introduced in B.2.2.
B.2.1 Image and Description Collection
The collection of images can significantly affect the effectiveness and robustness of the benchmark. To ensure the diversity and quality of images, we select 120 images from multiple commonly used benchmarks including MME [13], MMBench [36], MM-Vet [67], MMMU [70], MMHal-Bench [56], ScienceQA [37] and VCR [72].
Annotating a comprehensive description avoiding missing any important content in the image in a single turn can be overly challenging for even experienced annotators. In order to achieve reliable high content coverage and accuracy, we devise a three-step process as follows: (1) We first employ GPT-4V to generate detailed descriptions based on six different prompts listed in Table 8, where these prompts focus on different aspects of the same image. (2) We then merge these image descriptions, each with its own focus, to generate a draft comprehensive description by asking GPT-4 with prompt in Table 8. (2) Such a draft can be limited in both coverage and accuracy, so we ask human annotators to add more details and correct errors for each draft description. (3) To further ensure the comprehensiveness of image descriptions, each description is verified by at least two graduate students. (4) To ensure the accuracy and completeness of the annotations, the descriptions underwent a minimum of three rounds of additions and modifications. Specifically, the annotation price is 10 dollars per hour in average and only annotators with an English proficiency equivalent to a TOEFL score of 110 or higher are involved.
We provide three examples of annotated comprehensive image description and corresponding image in Figure 8, 9 and 10. Each image is paired with three different instructions.
B.2.2 Instruction Design
For each image, we design three related instructions to cover diverse scenarios. Specifically, inspired by MMBench [36] and MMMU [70], we focus on 8 important capabilities of MLLMs including:
  • Fine-grained perception refers to recognizing detailed aspects, such as characters, objects, and object attributes (e.g., color, material, shape).
  • Coarse perception primarily refers to the general visual content perception capability, which includes describing the image styles, atmosphere, scenes, etc.
  • Optical Character Recognition (OCR) involves the recognition of text and formulas in images.
  • Creative generation evaluates a model’s creative capabilities, including writing stories or advertisements derived from the image content, and critically analyzes the techniques of composition and photography.
  • Attribute reasoning primarily assesses the model’s capability to infer the style, subject, object function, person identity, and other aspects of images.
  • Relation reasoning primarily assesses the model’s capability to infer the relationships between different parts in the image, such as spatial relationships, inter-person relations, and other relationships among various elements.
  • Time series reasoning primarily assesses the capability to comprehend changes and predict future events across different scenarios depicted in an image.
  • Logical reasoning mainly assesses code comprehension and mathematical reasoning capabilities.
We present an example of an image with corresponding three instructions in Figure 8, where these instructions evaluate three different capabilities including fine-grained perception, relation reasoning, and coarse-grained perception.
To prevent over-fitting of the dataset, we randomly sampled the RefoMB dataset based on the proportions of each category, dividing it into dev and test splits. The dev split contain 99 images, while the test split comprises 261 images. Statistics of instructions in RefoMB are shown in Table 4 and Figure 6. In this paper, we initially release the dev split for MLLMs evaluation. The test split will be released after the dev split for six months.
B.3 Analytical Results
In this section, we analyze the reliability of RefoMB compared with other benchmarks and discuss the difference between using GPT-4 [44] or GPT-4V [43] as the evaluator.
B.3.1 Reliability Analysis of RefoMB
To explore the reliability of our evaluation results, we conduct an experiment comparing the human agreement of RefoMB with widely used MMHal-Bench [56], which all utilize GPT-4 as the evaluator. Specifically, we use these benchmarks to assess the performance of six commonly employed MLLMs, including LLaVA 1.5 [33], LLaVA-NeXT [35], GPT-4V [43], OmniLMM [46], RLAIF-7B and RLAIF-12B. For each instruction in every benchmark, we collect $302\times\binom{6}{2}=$ 30 response pairs by combining outputs generated by different models. Then, we uniformly sample 100 pairs from each benchmark and collect corresponding win-lose-tie decisions. For MMHal-Bench, which assigns absolute scores for each response, we compare the score value of two responses to get the decision. We then ask the human annotator to classify each evaluation result into “agree” or “disagree” and present the results in Table 5. We observe that RefoMB exhibits both higher reliability and more evenly distributed win and lose counts.
Benchmark Win/Lose/Tie Human Agree
MMHal Bench 38/28/34 85/100
RefoMB 45/46/ 9 96/100
Table 5: Human agreement of different hallucination-related benchmarks.
B.3.2 GPT-4 or GPT-4V as Evaluator
Compared with GPT-4 which handles text-only inputs, GPT-4V is specifically designed to handle multimodal inputs (text and visuals). Therefore, a natural question arises: Why not use GPT-4V as the evaluator which can directly perceive the image without relying on the image description? GPT-4V exhibits significant hallucination problems when perceiving images [66], which interferes with the reliability of evaluation, and we empirically find that GPT-4V always misunderstands the existence and number of objects, which agrees with [77]. To tackle these issues, we complete the perceiving process via an elaborately designed image description annotation process and ask GPT-4 to use the text-only description as an evaluation reference.
B.4 Example of Evaluation Results
To provide a more intuitive understanding of evaluation results on different tasks. As shown in Figure 11, we show a case of evaluation result from RefoMB.
B.5 RefoMB Dev Split Evaluation Results
We report the full evaluation results on the dev split of RefoMB in Table 6 including the trustworthiness win rate and overall win rate of each category.
B.6 RefoMB Test Split Evaluation Results
We report the full evaluation results on the test split of RefoMB in Table 7 including the trustworthiness win rate and overall win rate of each category.
Model Fine-grained Coarse Creative OCR Relation Attribute Logical Time series Average
Perception Perception Generation Reasoning Reasoning Reasoning Reasoning
Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win
VCD [25] 64.6 20.8 44.7 26.3 41.7 12.5 22.7 13.6 7.1 3.6 37.5 12.5 0.0 0.0 62.5 37.5 39.9 16.7
Less-is-more [71] 66.7 20.8 36.8 21.1 62.5 20.8 13.6 4.5 28.6 17.9 37.5 8.3 50.0 0.0 0.0 12.5 42.9 16.2
OPERA [21] 50.0 22.9 39.5 13.2 29.2 20.8 13.6 9.1 10.7 3.6 33.3 8.3 33.3 0.0 62.5 0.0 33.8 13.1
LURE [79] 45.8 6.2 31.6 5.3 25.0 0.0 18.2 4.5 17.9 0.0 12.5 0.0 33.3 0.0 62.5 0.0 29.8 3.0
Qwen-VL [4] 60.4 25.0 44.7 18.4 50.0 33.3 22.7 9.1 32.1 7.1 25.0 12.5 0.0 0.0 37.5 12.5 40.9 17.7
LLaVA-NeXT [35] 50.0 37.5 52.6 42.1 45.8 50.0 36.4 22.7 46.4 35.7 37.5 25.0 0.0 0.0 37.5 37.5 44.4 35.4
MiniGemini [31] 56.2 41.7 47.4 39.5 58.3 41.7 40.9 36.4 50.0 32.1 45.8 20.8 0.0 0.0 75.0 75.0 50.0 36.9
HA-DPO [76] 75.0 29.2 18.4 15.8 45.8 16.7 36.4 9.1 28.6 21.4 29.2 8.3 16.7 0.0 12.5 0.0 39.9 17.2
POVID [78] 58.3 22.9 52.6 18.4 62.5 20.8 4.5 4.5 32.1 7.1 50.0 4.2 0.0 0.0 37.5 0.0 44.4 13.6
LLaVA-RLHF [56] 39.6 18.8 36.8 26.3 37.5 25.0 13.6 4.5 7.1 14.3 12.5 8.3 0.0 0.0 25.0 25.0 26.3 17.2
Silkie [28] 60.4 29.2 28.9 26.3 45.8 33.3 22.7 13.6 32.1 10.7 37.5 12.5 0.0 0.0 37.5 12.5 38.9 21.2
RLHF-V [66] 50.0 22.9 52.6 28.9 20.8 4.2 36.4 4.5 32.1 14.3 45.8 29.2 50.0 0.0 25.0 0.0 41.4 17.7
LLaVA 1.5 [33] 54.9 20.1 40.4 18.4 34.7 23.6 15.2 4.6 33.3 13.1 29.2 11.1 0.0 0.0 41.7 29.2 36.9 16.2
+ RLAIF-V 68.2 27.6 51.3 30.9 51.0 30.2 26.1 18.2 42.0 19.6 38.5 11.5 16.7 0.0 15.6 0.0 47.2 22.5
+ RLAIF-V BoN 70.3 28.6 57.2 26.3 65.6 40.6 36.4 22.7 63.4 25.9 41.7 10.4 12.5 0.0 31.3 0.0 55.7 24.4
OmniLMM [46] 55.6 27.4 50.9 14.9 56.2 22.2 26.5 19.7 33.3 16.1 40.3 15.3 11.1 0.0 43.8 0.0 44.7 18.5
+ RLAIF-V 75.0 45.8 57.9 29.0 66.7 41.7 31.8 4.6 57.1 17.9 45.8 29.2 33.3 0.0 62.5 0.0 58.1 28.3
+ RLAIF-V BoN 84.2 34.6 62.6 42.6 75.8 34.2 35.5 10.0 57.1 25.0 50.8 33.3 20.0 0.0 62.5 22.5 62.9 30.3
GPT-4V [43] 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0
Table 6: The trustworthiness win rate / overall win rate of different MLLMs on eight capabilities of RefoMB dev split. Trust.: trustworthiness win rate, Win.: overall win-rate.
Model Fine-grained Coarse Creative OCR Relation Attribute Logical Time series Average
Perception Perception Generation Reasoning Reasoning Reasoning Reasoning
Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win Trust. Win
MiniGemini [31] 51.6 34.4 51.0 42.0 42.4 25.8 41.1 37.5 51.3 48.7 43.8 34.4 37.5 31.2 59.1 59.1 48.1 38.1
LLaVA 1.5 [33] 50.0 15.6 31.0 18.0 22.7 6.1 33.9 19.6 36.8 22.4 42.2 15.6 12.5 0.0 40.9 9.1 36.8 15.5
+ RLAIF-V 59.8 18.0 46.0 21.0 39.4 12.1 37.5 17.9 39.5 29.0 35.9 15.6 31.3 0.0 36.4 9.1 44.4 18.2
+ RLAIF-V BoN 66.4 20.5 51.0 25.0 47.0 12.1 35.7 16.1 38.2 25.0 37.5 18.8 37.5 0.0 54.6 0.0 48.7 18.8
OmniLMM [46] 54.1 15.6 56.0 25.0 43.9 6.1 33.9 14.3 35.5 25.0 48.4 17.2 6.3 0.0 36.4 0.0 45.4 16.5
+ RLAIF-V 65.6 26.5 55.0 29.7 54.0 18.7 32.1 16.7 56.6 39.5 55.7 25.0 29.2 6.3 63.6 21.2 54.8 25.9
+ RLAIF-V BoN 65.8 32.2 61.3 31.0 53.5 14.1 40.5 15.5 56.6 31.1 53.6 22.4 27.1 6.3 71.2 18.2 56.9 25.2
GPT-4V [43] 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0
Table 7: The trustworthiness win rate / overall win rate of different MLLMs on eight capabilities of RefoMB test split. Trust.: trustworthiness win rate, Win.: overall win-rate.
Prompts for Descriptions Collection Prompt for GPT4-V to Generate Image Descriptions:
As an expert in accurately and comprehensively describing visual information, you need to describe the components of an image as thoroughly and in as much detail as possible based on the questions provided. The generated description should enable a person who has not seen the image to reconstruct all its contents from your description alone. It is imperative that your answers are both accurate and comprehensive.
Principles:
The image description should be comprehensive while maintaining accuracy and avoid to introduce incorrect information that does not align with the image. Each question consists of several sub-questions that need to be answered. The image description should address all sub-questions without omission. The image description can include reasonable inferences based on the provided image information, but it should not deviate from the content expressed in the image. Appropriate justifications should be provided based on the content of the image. If the image contains mathematical problems, provide the answers along with the description of the problem. If the image contains code, describe the code text and provide its execution results. If the image contains high school-level knowledge (such as food chains or molecular models), use as professional language as possible to describe the knowledge contained in the image, rather than merely describing the image content. The generated image description should be at least 700 words in length. Question: {Instruction}
Instructions List: Please observe and describe the experience or feelings elicited by this picture, discussing aspects such as style, theme, setting, mood, and quality. Please describe the overall style of the image along with your viewing experience or feelings, and provide a detailed analysis of the main compositional elements in the image, including shape, position, color, and texture among other visual characteristics. Based on the image, describe the events depicted and speculate on possible causes and consequences; explain how the relationships between various elements in the image support your predictions. Carefully observe the image, provide a detailed description of the image content and background, and explain the scene as well as any notable aspects of the composition of its elements. Please list as comprehensively and in as much detail as possible all the components you observe in the image, describing the details of these components including shape, position, color, texture, and other visual features, and explain the connections between these components. Describe the overall style of the image, detailing all the aspects that you find impressive or interesting, and describe the emotional responses and viewing experiences it conveys to you. Prompt for Merging Different Responses:
You are a text information integration expert. Currently, there are two texts describing an image from different perspectives. Your task is to integrate the information from these texts to form a comprehensive and detailed description. You must retain as much of the valid information from both texts as possible. Please note that if the integrated text contains content that is inconsistent with the given descriptions, you will face severe penalties.
- Description 1: {description A}
- Description 2: {description B}
Table 8: Prompts for GPT-4V image descriptions collection.
Figure 7: Prompts of the evaluation of RefoMB.
Figure 8: Example of samples in the RefoMB benchmark including the reference description, instructions and corresponding categories.
Figure 9: Example of samples in the RefoMB benchmark including the reference description, instructions and corresponding categories.
Figure 10: Example of samples in the RefoMB benchmark including the reference description, instructions and corresponding categories.
Figure 11: Example of evaluation results from RefoMB.
Appendix C Implementation Details
In this section, we introduce more implementation details of the RLAIF-V framework and our experimental results.
Method Agreement. Object MHuman. AMBER
HalBench
Acc. Resp. $\downarrow$ Ment. $\downarrow$ Resp. $\downarrow$ Acc. F1
REJ-P 83.3 27.1 13.9 53.4 78.1 84.9
REJ-C 96.7 13.3 7.5 41.8 79.9 85.9
Table 9: Performance of different combine strategies. Agreement.: Human agreement of the constructed pairs, MHuman.: MHumanEval.
C.1 Different Combine Strategies
Besides scoring each response with the number of rejected claims (REJ-N), we also try to use percentage of rejection (REJ-P), which counts the number $n_{rej}$ of claims that have $p_{no}>p_{yes}$ and $S_{i}=\frac{n_{rej}}{m}$. Comparison results of different combination methods are shown in Table 9. We observe that REJ-C obtains better pairwise accuracy and achieves promising hallucination reduction on Object HalBench and MHumanEval.
C.2 Response split and question generation
We collect 2k examples for both the claim extraction and question conversion task from the open-source Llama 3 70B [41] to train a small Llama 3 8B [41] model for efficient split and conversion. The data collection and fine-tuning process costs 1.2h and 0.5h with an 8xA100 80G machine separately.
C.3 No divide-and-conquer Feedback Collection
We list the prompt we used to collect feedback from MLLMs with self-rewarding [69] in Table 13, where we directly ask the open-source MLLM to generate the holistic helpfulness and trustworthiness score of a response.
Appendix D Analysis on RLHF-V Dataset
D.1 Response Generation Model
RLHF-V [66] relies on human annotators identify and correct hallucinations, whereas RLAIF-V obtains feedback from open-source models without requiring human labor. The high cost of RLHF-V makes it challenging to provide correctional feedback for each model. We investigated the open-source dataset [1] of RLHF-V and found there are no responses generated by LLaVA 1.5 7B. We list the detailed proportions of responses generated by different MLLMs in the RLHF-V dataset in Table 10.
Model Proportion
Muffin [65] 38.7%
LLaVA 1.0 [34] 28.8%
Zephyr_MM 14.6%
InstructBLIP [10] 13.1%
Qwen-VL-Chat [4]  4.9%
Table 10: Proportions of responses generated by different MLLMs in the RLHF-V Dataset.
D.2 Hallucination Distribution
Upon reviewing the detailed evaluation of different MLLMs on MMHal Bench [56], as shown in Table 11, we found significant variation in the fine-grained hallucination score across models. Specifically, the correlations between Muffin [65] and LLaVA 1.5 7B, and between LLaVA 1.0 and LLaVA 1.5 7B, are even negative. Since the RLHF-V dataset primarily includes data from these two models, its effectiveness is significantly reduced due to the limited shared hallucination distribution.
Model Correlation Attribute Adversarial Comparison Counting Relation Environment Holistic Other
LLaVA 1.0 [34] -0.08 0.67 0.00 1.75 1.58 1.50 1.25 1.50 0.67
Muffin [65] -0.09 1.92 3.00 1.25 1.67 1.25 2.33 1.92 2.08
OmniLMM [46] 0.81 4.92 3.33 3.00 2.42 3.42 3.42 1.75 2.83
LLaVA 1.5 [33] 1.00 3.83 2.08 2.75 1.75 2.17 2.67 2.00 1.67
Table 11: Fine-grained hallucination scores of different MLLMs on MMHal Bench, and their correlation with LLaVA 1.5.
Appendix E Qualitative Results
We provide more qualitative results in this section to better reflect the effectiveness of our method, as shown in Figure 12 and Figure 13.
Figure 12: Qualitative results of RLAIF-V 7B and RLAIF-V 12B compared with GPT-4V. Correct answers and hallucinations are highlighted in color respectively.
Figure 13: Qualitative results of and RLAIF-V 12B compared with GPT-4V. Correct answers and hallucinations are highlighted in color respectively.
Appendix F Potential Impact and Limitations
Our RLAIF-V framework is designed for constructing high-quality AI feedback for multimodal large language models to better align with human preference, especially for improving trustworthiness in visual-language conversation. Unlike approaches that rely on proprietary MLLMs or human feedback, our approach enables open-source MLLMs to learn and improve from peer feedback. We hope RLAIF-V can facilitate teams in the community to make their MLLMs more trustworthy. There are also possible limitations of our RLAIF-V framework. The first one is that our method relies on training MLLMs, which may require certain costs. The second limitation is that though with marked improvement, RLAIF-V models still suffer from hallucination. It is worth exploring new methods to further improve model trustworthiness. Regarding social impacts, RLAIF-V might facilitate the usage of MLLMs and thus cause either positive or negative impacts of AI tools.
Table 12: Prompts for response split and claim conversion.
Prompts for Response Split and Claim Conversion Split Claims:
You are an expert in extracting facts from the given question-answer pair for an image. Your task is to extract and rewrite the facts mentioned in the question-answer pair into self-contained sentences. Exclude opinions or subjective statements.
[…]
Convert Claims into Questions:
You are an expert at modifying a given declarative sentence into a general question sentence. Your task is to modify the given declarative sentences one by one into a general question form. Do not change tenses or add extra content.
[…]
Prompts for Self-Rewarding Feedback Collection Hallucination:
Review the user’s question and the corresponding response using the additive 3-point scoring system (i.e., the possible scores are 0, 1, 2, 3 exclusively) described below.
Points are accumulated based on the satisfaction of each criterion: - Add 1 point if the response does not contain any objects that are not present in the given image. - Add another point if the attributes and position of each object mentioned in the response match the picture. - Award a third point if the relation between each mentioned objects mentioned in the response match the picture.
⟨user-question⟩ {question} ⟨/user-question⟩⟨response⟩ {answer} ⟨/response⟩
After examining the user’s instruction and the response:
- First, briefly justify your total score. - Then, give the score (0 or 1 or 2 or 3) in a single line without any other information.
Helpfulness:
Review the user’s question and the corresponding response using the additive 3-point scoring system (i.e., the possible scores are 0, 1, 2, 3 exclusively) described below.
Points are accumulated based on the satisfaction of each criterion: - Add 1 point if the response is relevant to the user’s inquiry and the given image. - Add another point if the response is detailed and answers the basic elements of the user’s question in a useful way. - Award a third point if the response addresses the user’s question directly and comprehensively, and is well-organized and helpful.
⟨user-question⟩ {question} ⟨/user-question⟩⟨response⟩ {answer} ⟨/response⟩
After examining the user’s instruction and the response:
- First, briefly justify your total score. - Then, give the score (0 or 1 or 2 or 3) in a single line without any other information.
Table 13: Prompts for no divide-and-conquer feedback collection.